How many milligrams of hydrocortisone powder should be used in compounding the following prescription?

Hydrocortisone 2.5% In Hydrophilic Ointment qs to 60gm

Sig: Apply as directed.
Answer #1

Weight of solute = \frac{2.5}{100}\times\frac{x}{60}

                          x = 1.5 gms

The weight of hydrocortisone powder in milligrams is 1500 mg (to convert grams to milligrams multiply with 1000)